# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ VII-G1 Federated learning-Byzantine attack
Different from the existing data poisoning attacks that affect the data collection integrity of the training datasets, Fang et al. [164] propose Byzantine attacks against federated learning which affect the integrity of the learning phase in the training process. The Byzantine attacks are formulated as optimization problems. The performance evaluation on four real-world datasets (i.e., Breast Cancer Wisconsin (Diagnostic), CHMNIST, Fashion-MNIST, and MNIST datasets) shows that the attacks against Byzantine-robust federated learning methods can substantially increase the error rates.
@@ VII-H1 Model Inversion attack
Model inversion attacks are a type of inference attack, where an adversary attempts to infer sensitive information about the training data, such as inputs or parameters, by analyzing the output of the trained model. Fredrikson et al. [260] demonstrated the feasibility of such attacks on machine learning models trained on sensitive data, such as medical records, by using optimization techniques to reconstruct input data that results in a given model output.
@@ VII-H2 Membership inference attack
A membership inference attack (MIA) is this kind of security vulnerability exposure where an adversary seeks to investigate whether an intended sample has been utilized for training the target ML model or not, on the basis of the model’s behavior and output. Shokri et al. [261] conducted a black-box MIA attack against ML with a binary NN classification task-based formalization with a shadow learning technique to identify members of the training set from non-members. Tang et al. [262] classified these attacks into two categories, namely, 1) direct attacks, which often make a single query to interrogate the targeted sample in a direct manner, and 2) indirect attacks, which often makes numerous queries to interrogate vicinity of the targeted sample to deduce membership.
@@ VII-I Highlights the impacts of vulnerabilities in Peer-to-Peer IoT learning
In this section, we have reviewed a variety of attacks that can significantly disrupt operations and compromise data in peer-to-peer IoT learning. Backdoor and poisoning attacks introduce malicious behavior into models or influence learning processes, while adversarial examples and evasion attacks can cause widespread model malfunctions or incorrect classifications. Combined attacks utilize multiple methods, making detection difficult. Sybil and byzantine attacks can disproportionately impact model learning or introduce unpredictability into the system. Dropping attacks disrupt the efficiency of the learning process. Inference, model extraction, and inversion attacks pose significant threats to data privacy by inferring sensitive information, copying proprietary models, or reconstructing original inputs from outputs. These vulnerabilities underscore the importance of strong security measures to protect data integrity, user privacy, and model robustness and accuracy.
@@ VIII Defense Methods Against Edge Learning Vulnerabilities
The previous section clearly highlighted the potential risks facing future 6G-IoT systems, as AI will be a key player in ensuring the proper functioning, optimizing, and safeguarding of these systems. However, in order to take full advantage of AI’s benefits while avoiding its associated security risks, the cybersecurity research community has made significant efforts to make this possible. In this section, we will attempt to shed light on the importance of such efforts in a comprehensive manner and provide a classification of defensive mechanisms against ML attacks, as presented in Figure 12. In Table XV, we provide a list of state-of-the-art methods proposed for securing machine learning systems. Table XVI presents the summary of defense methods against federated machine learning vulnerabilities.
@@
TABLE XVI: Summary of defense methods against edge learning vulnerabilities.
@@ A Bayesian Compromise Detection (BCD) algorithm to solve an optimization problem
Generates equivalent member/non-member feedback to alleviate black-box membership inference threats
@@ VIII-A Training Phase Defense Methods
We start with the most vulnerable stage of ML, namely the training phase. We classify the deference mechanisms in the training phase against five threat categories, namely poisoning, backdoor, byzantine, Sybil, and combined attacks, as shown in Figure 13.
@@ VIII-A1 Poisoning Attacks Defense Mechanisms
  • Bio-inspired: these defensive mechanisms are motivated by biological functions or structures. For example, Xie et al. [255] propose an AI-targeted local poisoning attacks defense mechanism, which is inspired by the biological immune system and implemented in a multiple-party training environment. The proposed system supports antigen identification, immune reaction, and immunological recall with an adversarial pipeline, with no restrictions concerning malicious clients’ numbers or the period of time they participate in the process. Also, it can make adaptive judgments regarding the aggregation weights of different local models. Detailed experimental outcomes across both image and text datasets, including MNIST, CIFAR-10, IMDB, and SST-5, with different neural networks, including CNN and LSTM, and under different types of poisoning attacks, such as local label flipping attacks. The results proved the merit of the proposed system with respect to model efficiency and resilience against poisoning attacks. Specifically, the accuracy reaches 95% even with 75% of the participating clients being compromised. Another Bio-inspired technique used for defending against poisoning attacks is the cognitive computing paradigm, which attempts to reproduce human cognitive processes within a computerized model. Qu et al. [244] propose a decentralized cognitive computing model (D2C) using a Blockchained FL framework for IIoT, specifically, for Industry 4.0 systems. The authors introduced blockchain FL into cognitive computing learning to enhance its privacy-preserving abilities in a fully decentralized manner. In addition, the authors introduced a modified Markov decision process (MDP) for performance optimization purposes. Evaluation of the CIFAR-10 dataset regarding three different aspects, global model accuracy, model convergence, and resistance to poisoning attacks, indicated the feasibility of the proposed system.
@@
  • Experience-based Learning: such defensive mechanisms incorporate practical, real-life experiences into the systems learning workflow. Wang et al. [245] propose a Deep Q-network (DQN)-based feature selection approach for cleaning multi-source data and mitigating poisoning attacks, which is influenced by RL in terms of an experience-based learning paradigm. Specifically, the problem is formulated as a contest in dynamic states among agents and the environment. To circumvent the problem of high computational complexity, the authors provided SS, a spatial search algorithm for quicker learning of the DQN agent. Performed simulations on Beijing PM 2.5, Colon Data, and MNIST datasets proved that the proposed approaches can successfully mitigate data poisoning attacks.
@@
  • Preprocessor-based: these defensive mechanisms check the authenticity of the data before using it for training. Jangseung et al. [246] present Jangseung, a preprocessor for minimizing the effects of poisoning attacks without affecting model performance. The system is specifically developed to protect SVMs models from adversarial perturbations through the use of anomaly detection engines. The approach involves the identification of whether a given data point is an outlier based on historical training data, and newly entered data elements would be flagged and rejected as either faulty or malicious. The performance evaluation step using the MNIST and VCI Wisconsin breast cancer datasets involved training two replicate models under the same adversarial points, except one is protected by the proposed system. The results demonstrated that the protected model outperformed the unprotected model in all best-case scenarios, with 96.2% versus 53.2% with the MNIST dataset, as well as 88.18% versus 75.51% with the VCI breast cancer Wisconsin dataset.
@@
  • Reputation-Awareness: such defensive mechanisms check the genuineness level of the clients before allowing them to participate in the model training. Chen et al. [256] propose RAPFDL, a dynamic asynchronous anti-poisoning FL framework. The main concept is to provide a dynamic reputation measurement system among clients participating in learning. Individual clients can collect points by committing to update their local model parameters and exchange the earned points. As a result, customers are incentivized to download more updates to earn more reward points. The level of fairness is measured by the correlation coefficient between the accuracy of the final model and the individual contributions of the parties. Clients use a differentially private GAN (DPGAN) to generate artificial private data samples, and all of the exchange updates are encrypted using the improved additive homomorphic encryption. Then, the updates are stored in the blockchain as unmodifiable records, which ensures both audibility and transparency. It is demonstrated by the experimental results on MNIST, Fashion-MNIST, and CIFAR-10 datasets that RAPFDL can achieve a 30% reduction in learning time and is strong in resistance to representative poisoning attacks.
@@
  • Dimensionality-reduction and Clustering: these defensive mechanisms are designed to facilitate the data verification process. Li et al. [247] present a Kernel Principal Component Analysis (KPCA) and K-mean-based defense strategy against FL-targeted data-poisoning attacks, and precisely, label-flipping attacks. The main reason for this combination is that KPCA, an algorithm for dimensionality reduction is able to handle non-linear data efficiently while coping well with linear data (e.g. skewed data). The KPCA algorithm is consequently effective in locating and filtering out malicious updates, as these updates contain unique features, while k-mean clustering is used to reduce noise.
@@
  • Knowledge Transfer: in these defensive mechanisms, the knowledge already acquired is used to sustain new training. Chan et al. [248] propose a transfer learning-based defense against label-flipping attacks. While different studies aim to exclude malicious data from the training process. The authors considered the problem of making full use of contaminated samples in training, in cases where the first approach cannot be achieved. The problem is expressed as transfer learning where the footprint of malicious samples is downplayed by retrieving only information similar to non-malicious samples in the affected datasets. In addition, a weight initialization technique is proposed which allocates a sample weight according to the appropriateness of its predicted cost provided by a classifier against poisoning attacks.
@@
  • Cryptography-based Authentication: these defensive mechanisms are designed to detect any possible manipulation of the previously verified data, which protects the data from tampering. Stokes et al. [249] present a cryptography-based authentication technique, called VAMP, to protect the integrity of training data. To provide the metadata for the media objects, the manifest is used. The main idea behind this concept involves protecting the data sets for training and validation, as well as safeguarding any existing software package utilized to train and evaluate the model. To accomplish the protection of the data sets, manifests for the training and validation datasets such as metadata and data bindings are created, to be either published on the VAMP service or integrated directly into the respective datasets. Then, both the training and evaluation packages are then uploaded to the service. The proposed system uses SHA2 (256 or 512) to create manifests, while their serialization can be performed using JSON for textual datasets or CBOR for binary implemented datasets.
@@
  • Historical Distance Detection: In such mechanisms, the objective is to inspect each client update before using it. Shi et al. [250] propose a defense mechanism based on the analysis of the statistical relationship of the Euclidean distance between clients’ models. This defense mechanism is constructed based on the following statements. The distance between honest clients’ models is not similar to the distance between harmful clients’ models. The reason that most distances are similar is that the amount of malicious clients is much lower than the number of honest clients. Based on this, the main solution proposed is to force the aggregation server to choose the most possible benign clients, which are established on the minimal sum of their distance from each other. Experiment evaluations on MNIST with different data distribution techniques (IID and non-IID) validated the feasibility of the proposed system.
@@ VIII-A2 Backdoor Attacks Defense Mechanisms
  • Federated Filters: In these mechanisms, the implementation takes place on the aggregation server side and it can be used for client monitoring and security attack filter distribution. Hou et al. [257] focused on proposing a defensive mechanism against backdoor attacks in IIoT-FL environments. The authors trained multiple backdoor filters along with various combinations of XAI models and classifiers on the server side in order to guarantee the identification of the backdoor entries. In addition, the authors proposed a blur-label-flipping strategy to sanitize the backdoor locations, which allows reclaiming of data availability. The results of experimental evaluations on MNIST and CIFAR10 datasets prove the validity of the system with an accuracy of up to 99% in recognizing backdoor samples.
  • Feedback-based: such mechanisms are designed to continuously monitor and provide authenticity feedback on participating clients. Andreina et al. [251] propose a feedback-based FL backdoor detection technique called BaFFLE, which makes use of several clients’ data for both training and model tampering discovery. The defense strategy consists of taking advantage of the existence of diverse datasets across clients by embedding a feedback loop within the FL workflow. To address the various challenges of running a distributed protocol in a Byzantine environment, the authors used an off-the-shelf procedure to locally (clients-side) benchmark and compare the updated model’s classification performance against the former one, while excluding updates that demonstrate unexpected behavior. The proposed system is evaluated using the CIFAR-10 and FEMNIST datasets, which results show 100% accuracy with less than 5% false positive rate.

### v2

@@ VII-G1 Federated learning-Byzantine attack
Federated learning in 6G networks decentralizes training across multiple devices, boosting privacy and efficiency. However, a Byzantine attack, where malicious nodes provide faulty data or computations, can degrade the model’s accuracy and compromise the collaborative learning process. Without robust defense mechanisms, 6G networks might suffer reduced performance, security vulnerabilities, and data integrity issues. Different from the existing data poisoning attacks that affect the data collection integrity of the training datasets, Fang et al. [185] propose Byzantine attacks against federated learning which affect the integrity of the learning phase in the training process. The Byzantine attacks are formulated as optimization problems. The performance evaluation on four real-world datasets (i.e., Breast Cancer Wisconsin (Diagnostic), CHMNIST, Fashion-MNIST, and MNIST datasets) shows that the attacks against Byzantine-robust federated learning methods can substantially increase the error rates.
@@ VII-H1 Model Inversion attack
Model inversion attacks in 6G could exploit the advanced machine learning frameworks, potentially revealing user data processed by these algorithms. As 6G networks rely heavily on AI for optimizing and personalizing services, such attacks could undermine user privacy. Without robust security measures, compromised data confidentiality might overshadow the envisioned benefits of 6G. Model inversion attacks are a type of inference attack where an adversary attempts to infer sensitive information about the training data, such as inputs or parameters, by analyzing the output of the trained model. Fredrikson et al. [287] demonstrated the feasibility of such attacks on machine learning models trained on sensitive data, such as medical records, by using optimization techniques to reconstruct input data that results in a given model output.
@@ VII-H2 Membership inference attack
Membership inference attacks target machine learning models to determine whether a specific data point was part of the training set. In the context of 6G, which promises enhanced machine learning integration and ultra-reliable low-latency communications, such attacks could compromise user privacy by deducing user-specific data used during network optimizations. A membership inference attack (MIA) is this kind of security vulnerability exposure where an adversary seeks to investigate whether an intended sample has been utilized for training the target ML model or not based on the model’s behavior and output. Shokri et al. [288] conducted a blue-box MIA attack against ML with a binary NN classification task-based formalization with a shadow learning technique to identify members of the training set from non-members. Tang et al. [289] classified these attacks into two categories, namely, 1) direct attacks, which often make a single query to interrogate the targeted sample directly, and 2) indirect attacks, which often makes numerous queries to interrogate vicinity of the targeted sample to deduce membership.
@@ VII-I Highlights the impacts of vulnerabilities in Peer-to-Peer IoT learning
In this section, we have reviewed a variety of attacks that can significantly disrupt operations and compromise data in peer-to-peer IoT learning. Backdoor and poisoning attacks introduce malicious behavior into models or influence learning processes, while adversarial examples and evasion attacks can cause widespread model malfunctions or incorrect classifications. Combined attacks utilize multiple methods, making detection difficult [107]. Sybil and byzantine attacks can disproportionately impact model learning or introduce unpredictability into the system. Dropping attacks disrupt the efficiency of the learning process. Inference, model extraction, and inversion attacks threaten data privacy by inferring sensitive information, copying proprietary models, or reconstructing original inputs from outputs. These vulnerabilities underscore the importance of strong security measures to protect data integrity, user privacy, and model robustness and accuracy.
TABLE XVII: Impact of ML Attacks on 6G Features.
ML Attack Type
URLLC
AI-Centric Networking
D2D Communication
Ubiquitous Sensing
RNN Backdoor
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Imperceptible Backdoor Pattern Attack
$\checkmark$
$\bigstar$
Distributed Backdoor Attack
$\bigstar$
$\checkmark$
Pixel-space Backdoor Attack
$\checkmark$
$\checkmark$
Clean-label Backdoor Attack
$\bigstar$
$\checkmark$
White-box Adversarial Attack
$\bigstar$
$\checkmark$
Blue-box Adversarial Attack
$\bigstar$
$\checkmark$
Grey-box Adversarial Attack
$\bigstar$
$\checkmark$
Synergetic Attack
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Federated Poisoning Attack
$\bigstar$
$\checkmark$
Intrusion Poisoning Attack
$\checkmark$
$\checkmark$
$\bigstar$
Label Flipping Attack
$\checkmark$
$\checkmark$
Poisonous Label Attack
$\checkmark$
Clean-label Poisoning Attack
$\bigstar$
$\checkmark$
Data Poisoning Attack
$\checkmark$
$\checkmark$
Generative Poisoning Attack
$\checkmark$
$\bigstar$
Sybil-based Poisoning Attack
$\checkmark$
Sybil-based Collusion Attack
$\checkmark$
Federated learning-Byzantine Attack
$\checkmark$
$\bigstar$
$\checkmark$
Model Inversion Attack
$\checkmark$
$\bigstar$
Membership Inference Attack
$\checkmark$
$\checkmark$
$\checkmark$" indicates that the attack significantly impacts or exploits the specific feature of 6G.
$\bigstar$" indicates a partial impact on the specific 6G feature.
An empty cell indicates that the attack isn’t directly related to the 6G feature.
VII-J Impact of ML Attacks on 6G Features
As the next frontier in wireless communications, 6G is set to bring a myriad of advancements, particularly in areas such as Ultra-Reliable and Low-Latency Communication (URLLC), AI-Centric Networking, Device-to-Device (D2D) Communication, and Ubiquitous Sensing & Data Collection. However, with these advancements come potential vulnerabilities, especially in the context of machine learning attacks.
Table XVII gives an in-depth look into how various ML attack types may exploit or be impacted by the key features of 6G. Notably, RNN Backdoor Attacks appear to exploit all the highlighted 6G features, underlining the multifaceted threats posed by this attack type. Moreover, the interplay between AI-Centric Networking and various adversarial attacks, including White-box, Blue-box, and Grey-box, emphasizes the potential risks of deeply embedding AI into 6G’s core functions.
Device-to-Device communication, allowing for direct device interaction without centralized routing, presents another vulnerable point, susceptible to Distributed Backdoor Attacks, Federated Poisoning Attacks, and others. This shows the importance of establishing robust security mechanisms for direct communication protocols in 6G networks.
Lastly, new vulnerabilities may arise with 6G’s potential to expand sensing technologies and become more context-aware through Ubiquitous Sensing. This is evident in its potential susceptibility to Pixel-space Backdoor Attacks, Label Flipping Attacks, and Data Poisoning Attacks. As 6G develops, understanding and mitigating these vulnerabilities becomes paramount to ensure a secure, reliable network.
@@ VIII Defense Methods Against Edge Learning Vulnerabilities
The previous section highlighted the potential risks facing future 6G-IoT systems, as AI will be a key player in ensuring these systems’ proper functioning, optimizing, and safeguarding. However, to take full advantage of AI’s benefits while avoiding its associated security risks, the cybersecurity research community has made significant efforts to make this possible. In this section, we will attempt to comprehensively shed light on the importance of such efforts and provide a classification of defensive mechanisms against ML attacks, as presented in Figure 14. Table XVI lists state-of-the-art methods proposed for securing machine learning systems. Table XVIII summarizes defense methods against federated machine learning vulnerabilities.
@@
TABLE XVIII: Summary of defense methods against edge learning vulnerabilities.
@@ A Bayesian Compromise Detection (BCD) algorithm to solve an optimization problem
Generates equivalent member/non-member feedback to alleviate blue-box membership inference threats
@@ VIII-A Training Phase Defense Methods
We start with the most vulnerable stage of ML, namely the training phase. We classify the deference mechanisms in the training phase against five threat categories: poisoning, backdoor, byzantine, Sybil, and combined attacks, as shown in Figure 15.
@@ VIII-A1 Poisoning Attacks Defense Mechanisms
  • Bio-inspired: These defensive mechanisms are motivated by biological functions or structures. For example, Xie et al. [282] propose an AI-targeted local poisoning attacks defense mechanism, which is inspired by the biological immune system and implemented in a multiple-party training environment. The proposed system supports antigen identification, immune reaction, and immunological recall with an adversarial pipeline, with no restrictions concerning malicious clients’ numbers or the period they participate in the process. Also, it can make adaptive judgments regarding the aggregation weights of different local models. Detailed experimental outcomes across image and text datasets, including MNIST, CIFAR-10, IMDB, and SST-5, with different neural networks, including CNN and LSTM, and under different poisoning attacks, such as local label flipping attacks. The results proved the merit of the proposed system concerning model efficiency and resilience against poisoning attacks. Specifically, the accuracy reaches 95% even with 75% of the participating clients being compromised. Another Bio-inspired technique used for defending against poisoning attacks is the cognitive computing paradigm, which attempts to reproduce human cognitive processes within a computerized model. Qu et al. [271] propose a decentralized cognitive computing model (D2C) using a Blockchained FL framework for IIoT, specifically, for Industry 4.0 systems. The authors introduced blockchain FL into cognitive computing learning to enhance its privacy-preserving abilities in a fully decentralized manner. In addition, the authors introduced a modified Markov decision process (MDP) for performance optimization purposes. Evaluation of the CIFAR-10 dataset regarding three different aspects, global model accuracy, model convergence, and resistance to poisoning attacks, indicated the feasibility of the proposed system.
@@
  • Experience-based Learning: such defensive mechanisms incorporate practical, real-life experiences into the systems learning workflow. Wang et al. [272] propose a Deep Q-network (DQN)-based feature selection approach for cleaning multi-source data and mitigating poisoning attacks, which RL influences in terms of an experience-based learning paradigm. Specifically, the problem is formulated as a contest in dynamic states among agents and the environment. To circumvent the problem of high computational complexity, the authors provided SS, a spatial search algorithm for quicker learning of the DQN agent. Performed simulations on Beijing PM 2.5, Colon Data, and MNIST datasets proved that the proposed approaches can successfully mitigate data poisoning attacks.
@@
  • Preprocessor-based: These defensive mechanisms check the authenticity of the data before using it for training. Jangseung et al. [273] present Jangseung, a preprocessor for minimizing the effects of poisoning attacks without affecting model performance. The system is specifically developed to protect SVMs models from adversarial perturbations through anomaly detection engines. The approach involves identifying whether a given data point is an outlier based on historical training data, and newly entered data elements would be flagged and rejected as either faulty or malicious. The performance evaluation step using the MNIST and VCI Wisconsin breast cancer datasets involved training two replicate models under the same adversarial points, except one is protected by the proposed system. The results demonstrated that the protected model outperformed the unprotected model in all best-case scenarios, with 96.2% versus 53.2% with the MNIST dataset, as well as 88.18% versus 75.51% with the VCI breast cancer Wisconsin dataset.
@@
  • Reputation-Awareness: such defensive mechanisms check the genuineness level of the clients before allowing them to participate in the model training. Chen et al. [283] propose RAPFDL, a dynamic asynchronous anti-poisoning FL framework. The main concept is to provide a dynamic reputation measurement system among clients participating in learning. Individual clients can collect points by committing to update their local model parameters and exchange the earned points. As a result, customers are incentivized to download more updates to earn more reward points. The correlation coefficient between the accuracy of the final model and the individual contributions of the parties measures the level of fairness. Clients use a differentially private GAN (DPGAN) to generate artificial private data samples, and all of the exchange updates are encrypted using the improved additive homomorphic encryption. Then, the updates are stored in the blockchain as unmodifiable records, which ensures both audibility and transparency. It is demonstrated by the experimental results on MNIST, Fashion-MNIST, and CIFAR-10 datasets that RAPFDL can achieve a 30% reduction in learning time and is strong in resistance to representative poisoning attacks.
@@
  • Dimensionality-reduction and Clustering: These defensive mechanisms facilitate data verification. Li et al. [274] present a Kernel Principal Component Analysis (KPCA) and K-mean-based defense strategy against FL-targeted data-poisoning attacks, and precisely, label-flipping attacks. The main reason for this combination is that KPCA, an algorithm for dimensionality reduction can handle non-linear data efficiently while coping well with linear data (e.g., skewed data). The KPCA algorithm is consequently effective in locating and filtering out malicious updates, as these updates contain unique features, while k-mean clustering is used to reduce noise.
@@
  • Knowledge Transfer: In these defensive mechanisms, the knowledge already acquired is used to sustain new training. Chan et al. [275] propose a transfer learning-based defense against label-flipping attacks. At the same time, different studies aim to exclude malicious data from the training process. The authors considered the problem of fully using contaminated samples in training in cases where the first approach cannot be achieved. The problem is expressed as transfer learning where the footprint of malicious samples is downplayed by retrieving only information similar to non-malicious samples in the affected datasets. In addition, a weight initialization technique is proposed which allocates a sample weight according to the appropriateness of its predicted cost provided by a classifier against poisoning attacks.
@@
  • Cryptography-based Authentication: These defensive mechanisms are designed to detect any possible manipulation of the previously verified data, which protects the data from tampering. Stokes et al. [276] present a cryptography-based authentication technique, called VAMP, to protect the integrity of training data. To provide the metadata for the media objects, the manifest is used. The main idea behind this concept involves protecting the data sets for training and validation and safeguarding any existing software package utilized to train and evaluate the model. To protect, the data sets manifests for the training and validation datasets such as metadata and data bindings are created to be either published on the VAMP service or integrated directly into the respective datasets. Then, both the training and evaluation packages are then uploaded to the service. The proposed system uses SHA2 (256 or 512) to create manifests, while their serialization can be performed using JSON for textual datasets or CBOR for binary implemented datasets.
@@
  • Historical Distance Detection: In such mechanisms, the objective is to inspect each client update before using it. Shi et al. [277] propose a defense mechanism based on analyzing the statistical relationship of the Euclidean distance between clients’ models. This defense mechanism is constructed based on the following statements. The distance between honest clients’ models is not similar to the distance between harmful clients’ models. Most distances are similar because the number of malicious clients is much lower than that of honest clients. Based on this, the primary solution proposed is to force the aggregation server to choose the possible benign clients, which are established on the minimal sum of their distance from each other. Experiment evaluations on MNIST with different data distribution techniques (IID and non-IID) validated the feasibility of the proposed system.
@@ VIII-A2 Backdoor Attacks Defense Mechanisms
  • Federated Filters: In these mechanisms, the implementation takes place on the aggregation server side and it can be used for client monitoring and security attack filter distribution. Hou et al. [284] focused on proposing a defensive mechanism against backdoor attacks in IIoT-FL environments. The authors trained multiple backdoor filters and various combinations of XAI models and classifiers on the server side to guarantee the identification of the backdoor entries. In addition, the authors proposed a blur-label-flipping strategy to sanitize the backdoor locations, allowing data availability reclaiming. The results of experimental evaluations on MNIST and CIFAR10 datasets prove the system’s validity with an accuracy of up to 99% in recognizing backdoor samples.
  • Feedback-based: Such mechanisms are designed to monitor and provide authenticity feedback on participating clients continuously. Andreina et al. [278] propose a feedback-based FL backdoor detection technique called BaFFLE, which uses several clients’ data for training and model tampering discovery. The defense strategy consists of taking advantage of the existence of diverse datasets across clients by embedding a feedback loop within the FL workflow. To address the various challenges of running a distributed protocol in a Byzantine environment, the authors used an off-the-shelf procedure to locally (clients-side) benchmark and compare the updated model’s classification performance against the former one while excluding updates that demonstrate unexpected behavior. The proposed system is evaluated using the CIFAR-10 and FEMNIST datasets, which results show 100% accuracy with less than 5% false positive rate.